# Simple Policy Optimization

### v4

@@ 3 Method
[…]
4 Experiments
We report results on the Atari 2600 environment (Bellemare et al., 2013; Machado et al., 2018). The Atari 2600 environment is a popular benchmark in the field of reinforcement learning. It consists of a set of video game environments designed for the Atari 2600 gaming console. These environments serve as testing grounds for developing and evaluating reinforcement learning algorithms and have been widely used because of its diverse set of games, ranging from simple classics like Pong to more complex titles like SpaceInvaders and Breakout. This diversity allows researchers to assess the generalization capabilities of reinforcement learning algorithms across different tasks and complexities. We preprocessed the Atari 2600 environment, for example, the input to the network is the stacking of the last four frames resized to $84\times 84\times 4$, reward clipping and so on (Mnih et al., 2013; Huang et al., 2022). In all experiments, we use the hyperparameters provided in the Appendix A unless otherwise specified. For each environment in Atari 2600, we compute and visualize the standard deviation of the algorithm’s performance across 3 separate runs with different random seeds.
Figure 4: We over-optimize the objectives of PPO and SPO for 1000 epochs in the early stages of training, where $d_{\max}$ is set to 0.01, 0.015, 0.02, 0.025 and 0.03.
Our experiments are divided into four parts. The first part compares the performance of several PPO variants and the SPO algorithm in the Atari 2600 environment, as well as the changes in KL divergence. In the second part, we increase the number of update epochs from 8 to 1000 in a certain training stage to over-optimize the respective objectives of PPO and SPO, and observe their optimization of surrogate objectives and the growth trend of KL divergence. In the third part, we try to increase the depth of the neural network and observe how the complexity of the network structure affects the performance of PPO and SPO in the training process. Finally, we conducted a sensitivity analysis on the hyperparameter $d_{\max}$ of the SPO algorithm.
@@ 4.1 Comparison to PPOs
For the sake of experimental fairness, the policy network and value network structures of both the SPO and PPO algorithms are entirely identical, see Appendix F. They share a common convolutional layer designed to compress high-dimensional image inputs into a low-dimensional representation of 512 dimensions. Subsequently, a fully connected layer is connected separately to generate outputs for the policy network (with the output dimensionality corresponding to the action space) and the value network (with a 1-dimensional output). Partial experimental results are shown in Figure 3, with more detailed results provided in the Appendix C. Here, PPO-PENALTY (PPO-P) and PPO-CLIP with an early stopping strategy (PPO-E) are two main variants introduced to limit the excessive increase of KL divergence of PPO during the training process. We maintain consistency in the corresponding KL divergence thresholds for these two variants and the threshold $d_{\max}$ for SPO, all set to 0.02. It can be seen that SPO nearly achieves significantly better sample efficiency than other PPO variants in most environments, ranking second only to SPO-E in a few cases, while maintaining the lowest KL divergence throughout the entire training process. In contrast, both PPO-C and PPO-P consistently exhibit significantly high KL divergence. However, PPO-E is also able to effectively limit KL divergence, trailing only slightly behind SPO. We also notice that SPO exhibits the highest policy entropy during training, which indicates that SPO has a better exploration capability in the environment. We speculate that this may be a by-product of limiting the KL divergence. In fact, we conducted a sensitivity analysis on the hyperparameter $d_{\max}$ in Subsection 4.4, and we found that a lower value of $d_{\max}$ leads to a higher policy entropy.
(a) Training curves of return.
(b) Training curves of average KL divergence.
(c) Training curves of policy entropy.
@@
4.2 Over Optimization
To further evaluate whether SPO can satisfy trust region constraints under extreme conditions, we select four environments and over-optimize the objectives of PPO and SPO with different $d_{\max}$ for 1000 epochs in the early stages of training, results are shown in Figure 4. Where SPO-$x$ represents the SPO algorithm with the maximum KL divergence $d_{\max}$ set to $x$, PPO-$x$ represents the PPO algorithm with the ratio clipping parameter $\epsilon$ set to $x$. We can see that even in the case of over-optimization, SPO still has a very stable KL divergence, which remains below the threshold, and the surrogate objective gradually increases as the constraints become weaker ($d_{\max}$ increases). In contrast, PPO leads to excessively high average KL divergence and has relatively poor optimization performance for surrogate objective. Therefore, to a certain extent, this indicates that SPO optimizes the surrogate objective within the trust region as much as possible, and performs better than PPO.
(b) Training curves of average KL divergence.
(c) Training curves of policy entropy.
4.3 Increasing Network Complexity
We are interested in understanding how the depth or complexity of the neural networks [^1] Two neural network structures are shown in the appendix F. affects the performance of both PPO and SPO algorithms in the training process, as well as their KL divergence. This subsection aims to answer the following questions: When the network structure becomes more complex, will the ability of the SPO algorithm to limit KL divergence disappear? Will the phenomenon of PPO’s inability to limit KL divergence become more evident? We did not make any changes to the hyperparameters or other settings, except for increasing the number of network layers. PPO-1 and SPO-1 represent the use of shallow networks, while PPO-2 and SPO-2 represent the use of deeper networks. The results are shown in Figure 5. As the network becomes deeper, we can see that the KL divergence of PPO during the training process becomes larger than that in the shallow networks, sometimes leading to a sharp decline in its performance. In some environments, PPO almost loses its ability to learn effectively (e.g., Breakout and DemonAttack). In contrast, SPO demonstrates strong robustness to changes in network structure. Furthermore, we observed that the average KL divergence curves of SPO-1 and SPO-2 nearly overlap throughout the entire training process, which once again validates the effectiveness of SPO in limiting KL divergence. Additionally, this suggests that SPO is more likely to play a pivotal role in complex, large-scale neural network models, such as the Transformers (Vaswani et al., 2017) or DM (Diffusion Models) (Ho et al., 2020).
4.4 Sensitivity Analysis
We conducted a sensitivity analysis on the hyperparameter $d_{\max}$ of SPO, setting it to 0.01, 0.02, and 0.03, respectively. The experimental results are shown in Figure 6. We can see that the performance of the SPO algorithm is not highly sensitive to the selection of the hyperparameter $d_{\max}$, and the average KL divergence during the training process gradually increases as $d_{\max}$ increases. Notably, as $d_{\max}$ gradually decreases, the policy entropy of the SPO algorithm gradually increases. This result seems to be expected because the KL divergence $d$ satisfies
$\begin{split}d&=\sum_{a\in\mathcal{A}}\pi_{\theta_{\rm old}}(a|s_{t})\cdot\log\frac{\pi_{\theta_{\rm old}}(a|s_{t})}{\pi_{\theta}(a|s_{t})},\\ &=\underbrace{H\left(\pi_{\theta_{\text{old}}}(\cdot|s_{t}),\pi_{\theta}(\cdot|s_{t})\right)}_{\text{Cross Entropy}}-\underbrace{H\left(\pi_{\theta_{\text{old}}}(\cdot|s_{t})\right)}_{\text{Policy Entropy}}.\\ \end{split}$ (16)
The cross entropy is non-negative, and it will be a good choice for the agent to reduce the KL divergence by maintaining a higher policy entropy in the training process. We believe that there is a fundamental difference between SPO and PPO-E, the latter of which introduces an early stopping strategy to forcibly limit the KL divergence, as SPO always exhibits a higher policy entropy than PPO-E when they have a similar level of KL divergence.
@@ 5 Discussion
We propose Simple Policy Optimization (SPO), a first-order algorithm for policy optimization in reinforcement learning. Experimental results in the Atari 2600 environment demonstrate that compared to the traditional PPO algorithm, SPO significantly improves sample efficiency and effectively limits the average KL divergence between the old and current policies, while exhibiting higher policy entropy. SPO strikes a favorable balance between sample complexity, simplicity and wall-time. In this work, we do not advocate the use of the surrogate objective with ratio clipping. We show that the ratio clipping method is not effective in limiting the average KL divergence between the old and current policies, and this phenomenon becomes more pronounced as the network deepens. In contrast, SPO achieves higher sample efficiency with lower KL divergence and exhibits robustness to the depth or complexity of the network. We have noticed that some techniques employ reinforcement learning to allow LLMs (Large Language Models) to learn human preferences (Ouyang et al., 2022; Ziegler et al., 2019; Stiennon et al., 2020; Lee et al., 2023), called RLHF (Reinforcement Learning from Human Feedback). A crucial step in this process is to fine-tune the LLMs using policy optimization algorithms such as PPO. To prevent the “catastrophic forgetting” of LLMs, efforts have been focused on developing new algorithms that minimize the KL divergence with the base reference model during the fine-tuning process (Ziegler et al., 2019; Rafailov et al., 2024; Tunstall et al., 2023). It would be intriguing to apply SPO to this process, as the key difference lies in the fact that the old policy is no longer the policy network at the end of the previous training phase, but instead refers to the base reference policy.
@@ Appendix A Hyperparameters Setting
Table 1: Detailed hyperparameters
@@ B.1 Training Curves of Return
Figure 7: Training curves of return of PPO and SPO in Atari 2600 environments.
@@ Appendix C Comparison to PPOs
Appendix D Increasing Network Complexity
(a) Training curves of average KL divergence.
(b) Training curves of policy entropy.
Appendix E Sensitivity Analysis

### v5

@@ 4.1 Comparison to PPOs
For the sake of experimental fairness, the policy network and value network structures of both the SPO and PPO algorithms are entirely identical, see shallow network structure in Figure 4. They share a common convolutional layer designed to compress high-dimensional image inputs into a low-dimensional representation of 512 dimensions. Subsequently, a fully connected layer is connected separately to generate outputs for the policy network (with the output dimension corresponding to the action space) and the value network (with a 1-dimensional output).
@@
Partial experimental results are shown in Figure 2, with more detailed results provided in the Appendix B.4. Here, PPO-Penalty (PPO-P) and PPO-Clip with an early stopping strategy (PPO-E) are two main variants introduced to limit the excessive increase of KL divergence of PPO during the training process. We maintain consistency in the corresponding KL divergence thresholds for these two variants and the threshold $d_{\max}$ for SPO, all set to 0.02. It can be seen that SPO nearly achieves significantly better sample efficiency than other PPO variants in most environments, ranking second only to SPO-E in a few cases, while maintaining the lowest KL divergence throughout the entire training process (see Appendix B.4). In contrast, both PPO-C and PPO-P consistently exhibit significantly high KL divergence. However, PPO-E is also able to effectively limit KL divergence, trailing only slightly behind SPO.
We also notice that SPO exhibits the highest policy entropy during training (see Appendix B.4), which indicates that SPO has a better exploration of the environment. This reveals the fundamental difference between SPO and PPO-E, and we speculate that this may be a by-product of spontaneously limiting the KL divergence (not forcibly). In fact, we conducted a sensitivity analysis on the hyperparameter $d_{\max}$ in Appendix B.7, and we found that a lower value of $d_{\max}$ leads to a higher policy entropy.
Figure 3: We over-optimize the objectives of PPO and SPO for 1000 epochs in the early stages of training, where $d_{\max}$ is set to 0.01, 0.015, 0.02, 0.025 and 0.03, respectively.
4.2 Over Optimization
To further evaluate whether SPO can satisfy trust region constraints under extreme conditions, we select three environments and over-optimize the objectives of PPO and SPO with different $d_{\max}$ for 1000 epochs in the early stages of training, results are shown in Figure 3. Where SPO-$x$ represents the SPO algorithm with the maximum KL divergence $d_{\max}$ set to $x$, PPO-$x$ represents the PPO algorithm with the ratio clipping parameter $\epsilon$ set to $x$. We can see that even in the case of over-optimization, SPO still has a very stable KL divergence, which remains below the threshold, and the surrogate objective gradually increases as the constraints become weaker ($d_{\max}$ increases). In contrast, PPO always leads to excessively high average KL divergence and has relatively poor optimization performance for surrogate objective. Therefore, to a certain extent, this indicates that SPO optimizes the surrogate objective within the trust region as much as possible, and performs better than PPO.
Figure 4: Shallow network and deeper network structure in Atari 2600 environments.
4.3 Increasing Network Complexity
We are interested in understanding how the depth or complexity of the neural networks affects the performance of both PPO and SPO algorithms in the training process, as well as their KL divergence. This subsection aims to answer the following questions: When the network structure becomes more complex, will the ability of the SPO algorithm to limit KL divergence disappear? Will the phenomenon of PPO’s inability to limit KL divergence become more evident? The Two structures are shown in Figure 4.
We did not make any changes to the hyperparameters or other settings, except for increasing the number of network layers. PPO-1 and SPO-1 represent the use of shallow networks, while PPO-2 and SPO-2 represent the use of deeper networks. The results are shown in Figure 5. As the network becomes deeper, we can see that the KL divergence of PPO during the training process becomes larger than that in the shallow networks, sometimes leading to a sharp decline in its performance. In some environments, PPO almost loses its ability to learn effectively (e.g., Breakout and DemonAttack). In contrast, SPO demonstrates strong robustness to changes in network structure. Furthermore, we observed that the average KL divergence curves of SPO-1 and SPO-2 nearly overlap throughout the entire training process, which once again validates the effectiveness of SPO in limiting KL divergence. Additionally, this suggests that SPO is more likely to play a pivotal role in complex, large-scale neural network models, such as the Transformers [34] or DMs (Diffusion Models) [35].
@@ 4.3 Increasing Network Complexity
Figure 6: Training curves in MuJoCo environments.
4.4 The Scalability of SPO for Continuous Action Space
We further tested the scalability of the SPO algorithm in a continuous action space environment, e.g., MuJoCo [33], and surprisingly, SPO still performed exceptionally well.
Typically, in the continuous action space, we assume that the output of the policy network follows a multivariate Gaussian distribution, and we further assume that the components are mutually independent. Therefore, the KL divergence between them is the sum of the KL divergences of the individual components.
Now we assume that two random variables follow Gaussian distributions: $p(x)=\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and $q(x)=\mathcal{N}(\mu_{2},\sigma_{2}^{2})$, then their KL divergence is
$D_{\text{KL}}(p,q)=\log\frac{\sigma_{2}}{\sigma_{1}}+\frac{(\mu_{1}-\mu_{2})^{2}+\sigma_{1}^{2}-\sigma_{2}^{2}}{2\sigma_{2}^{2}}.$ (19)
The proof see Appendix B.6. For each environment in MuJoCo, we compute and visualize the standard deviation of the algorithm’s performance across 5 separate runs with different random seeds, the results are shown in Figure 6, Table 1 and Table 2.
Table 1: Average return of the last 10% training steps in MuJoCo environments.
Algorithm Ant-v4 HalfCheetah-v4 Hopper-v4 Humanoid-v4 HumanoidStandup-v4 Walker2d-v4
SPO-0.005 $4775.95^{\downarrow 440}$ $1623.82^{\color[rgb]{1,0,0}\uparrow 16}$ $1434.74^{\downarrow 37}$ $3587.51^{\color[rgb]{1,0,0}\uparrow 2765}$ $159441.18^{\color[rgb]{1,0,0}\uparrow 8996}$ $3187.8^{\downarrow 30}$
SPO-0.01 $5066.14^{\downarrow 150}$ $2328.6^{\color[rgb]{1,0,0}\uparrow 721}$ $1377.1^{\downarrow 95}$ ${\bf 3772.78}^{\color[rgb]{1,0,0}\uparrow 2950}$ $158407.45^{\color[rgb]{1,0,0}\uparrow 7962}$ ${\bf 3508.62}^{\color[rgb]{1,0,0}\uparrow 291}$
SPO-0.03 $4628.01^{\downarrow 588}$ ${\bf 2619.4}^{\color[rgb]{1,0,0}\uparrow 1012}$ ${\bf 1632.63}^{\color[rgb]{1,0,0}\uparrow 161}$ $3168.32^{\color[rgb]{1,0,0}\uparrow 2346}$ ${\bf 159841.75}^{\color[rgb]{1,0,0}\uparrow 9396}$ $3286.5^{\color[rgb]{1,0,0}\uparrow 69}$
PPO 5216.36 1607.55 1471.81 822.57 150445.3 3217.51
Table 2: Average return throughout the entire training process in MuJoCo environments.
Algorithm Ant-v4 HalfCheetah-v4 Hopper-v4 Humanoid-v4 HumanoidStandup-v4 Walker2d-v4
SPO-0.005 $1466.71^{\downarrow 524}$ $1229.05^{\downarrow 122}$ $1046.8^{\downarrow 109}$ $1318.69^{\color[rgb]{1,0,0}\uparrow 525}$ $132000.34^{\downarrow 4209}$ $1530.99^{\downarrow 362}$
SPO-0.01 $1249.85^{\downarrow 741}$ $1705.9^{\color[rgb]{1,0,0}\uparrow 355}$ $1088.38^{\downarrow 67}$ $1457.89^{\color[rgb]{1,0,0}\uparrow 664}$ $136563.02^{\color[rgb]{1,0,0}\uparrow 354}$ $1803.99^{\downarrow 89}$
SPO-0.03 $1035.62^{\downarrow 955}$ ${\bf 1861.74}^{\color[rgb]{1,0,0}\uparrow 511}$ $1132.27^{\downarrow 23}$ ${\bf 1501.02}^{\color[rgb]{1,0,0}\uparrow 708}$ ${\bf 136769.98}^{\color[rgb]{1,0,0}\uparrow 561}$ $1788.93^{\downarrow 104}$
PPO 1991.05 1351.08 1155.63 793.51 136209.27 1893.36
From Figure 6, we can see that the sample efficiency of SPO in the Humanoid-v4 environment is much better than PPO. This may be because the environment is relatively complex, and the KL divergence of PPO becomes too large (see Appendix B.6).
As we can see from Table 1, the SPO algorithm performs better than PPO in the last 10% training steps, especially for SPO-0.03, which outperforms PPO in all environments except Ant-v4. However, according to Table 2, the average performance of the SPO algorithm decreases throughout the entire training process. Overall, this indicates that SPO is capable of achieving or even surpassing the performance of PPO in continuous control tasks (especially tasks like Humanoid-v4, which are relatively more complex), because we value the final performance of the algorithm more than the average performance of the entire training process.
@@ 5 Discussion
We propose Simple Policy Optimization (SPO), a first-order algorithm for policy optimization in reinforcement learning. Experimental results in the Atari 2600 environment demonstrate that compared to the traditional PPO algorithm, SPO significantly improves sample efficiency and effectively limits the average KL divergence between the old and current policies, while exhibiting higher policy entropy. SPO strikes a favorable balance between sample complexity, simplicity and wall-time.
In this work, we do not advocate the use of the surrogate objective with ratio clipping. We show that the ratio clipping method is not effective in limiting the average KL divergence between the old and current policies, and this phenomenon becomes more pronounced as the network deepens. In contrast, SPO achieves higher sample efficiency with lower KL divergence and exhibits robustness to the depth or complexity of the network.
We have noticed that some techniques employ reinforcement learning to allow LLMs (Large Language Models) to learn human preferences [36, 37, 38, 39], called RLHF (Reinforcement Learning from Human Feedback). A crucial step in this process is to fine-tune the LLMs using policy optimization algorithms such as PPO. To prevent the “catastrophic forgetting” of LLMs, efforts have been focused on developing new algorithms that minimize the KL divergence with the base reference model during the fine-tuning process [37, 40, 41]. It would be intriguing to apply SPO to this process, as the key difference lies in the fact that the old policy is no longer the policy network at the end of the previous training phase, but instead refers to the base reference policy.
@@ Appendix A Hyperparameters Setting
A.1 Hyperparameters Setting in Atari 2600
Table 3: Detailed hyperparameters in Atari 2600 environments.
@@ A.1 Hyperparameters Setting in Atari 2600
A.2 Hyperparameters Setting in MuJoCo
Table 4: Detailed hyperparameters in MuJoCo environments.
Algorithm PPO SPO
Number of actors 8 8
Horizon 256 256
Learning rate $3\times 10^{-4}$ $3\times 10^{-4}$
Learning rate decay Linear Linear
Optimizer Adam Adam
Total steps $3\times 10^{6}$, $1\times 10^{7}$ $3\times 10^{6}$, $1\times 10^{7}$
Batch size 2048 2048
Update epochs 10 10
Mini-batch size 512 512
Mini-batches 4 4
GAE parameter $\lambda$ 0.95 0.95
Discount factor $\gamma$ 0.99 0.99
[…]
Value loss coeff. $c_{1}$ 0.5 0.5
Entropy loss coeff. $c_{2}$ 0.0 0.0
[…]
@@ B.1 Training Curves of Return
Figure 7: Training curves of return in Atari 2600 environments.
@@ B.4 Comparison to PPOs
B.5 Increasing Network Complexity